OpenStudy (anonymous):

1+cos(18x)=____. a) 2cos^2 (18x) b) 2sin^2 (18x) c) 2cos^2 (9x) d) 2sin^2 (9x)

OpenStudy (anonymous):

@Agl202 @jim_thompson5910 @Luigi0210

OpenStudy (anonymous):

@Luigi0210 plz help

OpenStudy (luigi0210):

Is it only \(1+\cos(18x) \)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats the whole question

OpenStudy (anonymous):

Isn't it C?

OpenStudy (luigi0210):

It would be, yes :) http://prntscr.com/7e0cn4

OpenStudy (anonymous):

THANK YOU!!

OpenStudy (anonymous):

\[1+\cos 2\theta=2\cos ^2\theta\]

OpenStudy (luigi0210):

No problem :)